4) Solve for x(t)n using Laplace transform: dx dt + 2x = 6 where x(0)=-5 Highlight the transient responce
Answer #1

dt de +22=6 By applying Islace transform SXC) –xco) +2X(s) = 5 Given, X(0)=5 XC5) (5+2)-S = XCS) = 6 +5 S 5(5+2) sta X (s) =